### v2

# Demystifying Multilingual Chain-of-Thought in Process Reward Modeling

Abstract
Large language models (LLMs) are designed to perform a wide range of tasks. To improve their ability to solve complex problems requiring multi-step reasoning, recent research leverages process reward modeling to provide fine-grained feedback at each step of the reasoning process for reinforcement learning (RL), but it predominantly focuses on English. In this paper, we tackle the critical challenge of extending process reward models (PRMs) to multilingual settings. To achieve this, we train multilingual PRMs on a dataset spanning seven languages, which is translated from English. Through comprehensive evaluations on two widely used reasoning benchmarks across 11 languages, we demonstrate that multilingual PRMs not only improve average accuracy but also reduce early-stage reasoning errors. Furthermore, our results highlight the sensitivity of multilingual PRMs to both the number of training languages and the volume of English data, while also uncovering the benefits arising from more candidate responses and trainable parameters. This work opens promising avenues for robust multilingual applications in complex, multi-step reasoning tasks. In addition, we release the code to foster research along this line. [^1] https://github.com/weixuan-wang123/Multilingual-PRM
Demystifying Multilingual Chain-of-Thought in Process Reward Modeling
Weixuan Wang1   Minghao Wu2   Barry Haddow1   Alexandra Birch1 1School of Informatics, University of Edinburgh 2Monash University {weixuan.wang, bhaddow, a.birch}@ed.ac.uk minghao.wu@monash.edu
1 Introduction
Aligning large language models (LLMs) with human preferences can significantly improve the model performance across various downstream tasks (Christiano et al., 2017; Ziegler et al., 2019). This requires a reward model that is trained on human preference data (Ziegler et al., 2019; Stiennon et al., 2020; Shen et al., 2021; Ouyang et al., 2022). Typically, reward models are trained based on the final outcome of the LLMs’ response, and we refer to these as outcome reward models (ORMs) (Cobbe et al., 2021a; Uesato et al., 2022; Yu et al., 2023a). However, most of recent work demonstrates that ORMs fall short on complex multi-step reasoning tasks (Uesato et al., 2022; Shao et al., 2024). To overcome this limitation, process reward models (PRMs) are introduced, providing fine-grained rewards at each step of the LLMs’ reasoning (Lightman et al., 2024; Li et al., 2023; Wang et al., 2024b; Ma et al., 2023). Previous research has shown that LLMs supervised by PRMs can effectively produce better responses (Wang et al., 2024b; Shao et al., 2024).
Despite these significant advances, recent research on ORMs and PRMs has predominantly focused on monolingual settings, particularly English (Lightman et al., 2024; Wang et al., 2024a, b). However, the exploration of multilingual PRMs remains relatively limited. Therefore, with the advent of multilingual LLMs, a natural research question arises: How can we effectively train multilingual PRMs for complex, multi-step reasoning tasks?
To address this research question, we translate the existing PRM datasets, PRM800K (Lightman et al., 2024) and Math-Shepherd (Wang et al., 2024b), from English into six additional languages, resulting in a total of seven seen languages for training. We then train multilingual PRMs using the collection of these translated datasets. We define three PRM setups: PRM-mono, PRM-cross, and PRM-multi. The PRM-mono setup is trained and evaluated solely on a single language, the PRM-cross setup is trained on one language but evaluated on all test languages, and the PRM-multi setup is trained on seven seen languages and evaluated on all test languages. Finally, we conduct a comprehensive evaluation on two popular reasoning tasks (MATH500 and MGSM) across 11 languages (seven seen languages and four unseen languages) using three LLMs (MetaMath-Mistral-7B, Llama-3.1-8B-math, and DeepSeekMath-7B-Instruct).
Our main takeaways are summarized as follows:
  • Multilingual PRM consistently outperforms monolingual and cross-lingual PRMs across all three LLMs. Our results demonstrate that PRM-multi significantly improves model performance, boosting average accuracy by up to +1.2 and +1.5 points compared to PRM-cross and PRM-mono, respectively (see Section 5.1).
  • Multilingual PRM is sensitive to both the number of languages and the amount of English training data. Our experiment shows that training an optimal multilingual PRM requires careful consideration of how many languages to include (see Section 5.2) and how much English data to use (see Section 5.3).
  • Multilingual PRM produces fewer errors in the early steps. We identify the first occurrences of wrong predictions made by PRMs and observe that PRM-multi produces fewer errors in the early steps compared to PRM-mono and PRM-cross (see Section 6.1).
  • Multilingual PRM can benefit more from more candidate responses and trainable parameters. Our analysis demonstrates that PRM-multi becomes more advantageous with a larger number of candidate responses (see Section 6.2) and when more trainable parameters are introduced (see Section 6.3).
2 Related Work
Reward Model in Mathematical Reasoning
To advance the accuracy of mathematical reasoning, reward models (RMs) have emerged as powerful tools for evaluating and guiding solution generation. In particular, two principal RM paradigms have garnered significant attention: the Outcome Reward Models (ORMs) (Cobbe et al., 2021a; Yu et al., 2023a) and the Process Reward Models (PRMs) (Uesato et al., 2022; Lightman et al., 2024; Li et al., 2023; Ma et al., 2023; Wang et al., 2024b; Luo et al., 2024; Gao et al., 2024; Wang et al., 2024a). ORMs assign a single score to an entire solution and thereby focuses on final correctness, whereas PRMs score each individual step of the reasoning process, offering more finer-grained evaluations. As a result, PRMs provide more detailed guidance and have demonstrated greater potential in enhancing reasoning capabilities compared to ORMs (Lightman et al., 2024; Wu et al., 2023).
Multilingual Reward Model
Beyond English-language tasks, the integration of RMs into multilingual scenarios is still under-explored. Reinforcement learning approaches often rely on RMs predominantly trained on English data (Shao et al., 2024; Yang et al., 2024a). This over-representation introduces biases, as these RMs may overfit to English-specific syntactic and semantic patterns, limiting their effectiveness in cross-lingual tasks and motivating the development of multilingual RMs (Hong et al., 2024). While there is growing evidence that cross-lingual transfer is feasible (Wu et al., 2024a; Hong et al., 2024), existing research often overlooks the unique challenges of multilingual reasoning. After the release of the OpenAI-o1 model (OpenAI, 2024), PRMs, with their capability for fine-grained feedback, have attracted even greater interest. Yet, the performance of multilingual PRMs in diverse linguistic contexts remains insufficiently investigated (Yang et al., 2024b). To bridge this gap, we investigate how multilingual PRMs contribute to solving mathematical tasks across different languages, aiming to provide insights into how fine-grained process supervision can enhance reasoning capabilities beyond English, thereby contributing to the development of more universally applicable reasoning models.
Figure 1: Framework of PRM.
3 Process Reward Modeling
3.1 PRM Training
Given a question $p$ and its solution $s$, the ORM assigns a single value to $s$ to indicate if $s$ is correct. We stack a binary classifier on top of the LLM and train the ORM with the binary cross-entropy loss:
$\mathcal{L}_{\text{ORM}}$ $=$ (1)
$-{(y_{s}\log(r_{s})+(1-y_{s})\log(1-r_{s}))}$
where $y_{s}$ is the ground truth label for the solution $s$ ($y_{s}=1$ if $s$ is correct, otherwise $y_{s}=0$), and $r_{s}$ is the probability score that $s$ is correct.
In contrast, the PRM evaluates each reasoning step of the solution $s$. The PRM is trained using the following loss function:
$\mathcal{L}_{\text{PRM}}$ $=$ (2)
$-\sum_{i=1}^{K}$ $y_{s_{i}}\log(r_{s_{i}})+(1-y_{s_{i}})\log(1-r_{s_{i}})$
where $s_{i}$ is the $i$-th step of the solution $s$, $y_{s_{i}}$ is the ground truth label for step $s_{i}$, $r_{s_{i}}$ is the score assigned to $s_{i}$ by the PRM, and $K$ is the total number of reasoning steps in the solution $s$. Compared to ORM, PRM provides more detailed and reliable feedback by evaluating individual steps.
3.2 Ranking for Verification
Following Wu et al. (2024b); Lightman et al. (2024); Wang et al. (2024b), we evaluate the performance of PRM using the best-of-N selection evaluation paradigm (Charniak and Johnson, 2005; Cobbe et al., 2021b). Specifically, given a question, multiple solutions are sampled from an LLM (referred to as the generator) and re-ranked using a reward model (referred to as the verifier). For each solution, as shown in Figure 1, PRM assesses the correctness of each reasoning step. The scores for all steps are averaged to compute an overall score for the solution. The highest-scoring solution is then selected as the final output. This approach enhances the likelihood of selecting solutions containing correct answers, thereby improving the success rate of solving mathematical problems with LLMs.
3.3 Reinforcement Learning with Process Supervision
Using the trained PRM, we fine-tune LLMs with Policy Optimization (PPO) (Schulman et al., 2017) in a step-by-step manner. This method differs from the conventional strategy that uses PPO with an ORM, that only gives a reward at the end of the response. Conversely, our step-by-step PPO offers rewards at the end of each reasoning step.
While we analyse PRM both intrinsically (using best-of-N), and extrinsically (using PPO), we focus on best-of-N for a clean testbed without confounders from reinforcement learning.
4 Experimental Setups
Training Datasets
We combine the PRM800K (Lightman et al., 2024) and Math-Shepherd (Wang et al., 2024b) as training data to finetune PRMs, and translate the combined dataset from English (en) to six languages: German (de), Spanish (es), French (fr), Russian (ru), Swahili (sw), and Chinese (zh) with using NLLB 3.3B (Costa-jussà et al., 2022). The reasoning step statistics are presented in Table 4 (Appendix A), and the parallel examples across seven languages have the same number of reasoning steps. To ensure high translation quality, we use regular expressions to filter out translated training instances that contain discrepancies in numbers or equations compared to the original English dataset.
Test Dataset
We evaluate the performance of LLMs using two widely used math reasoning datasets, MGSM (Shi et al., 2022) and MATH500 (Wang et al., 2024b). For the MATH500 datset, we translate it from English to ten languages: Bengali (bn), German (de), Spanish (es), French (fr), Japanese (ja), Russian (ru), Swahili (sw), Telugu (te), Thai (th), and Chinese (zh) with Google Translate, which is consistent with the languages included in the MGSM dataset. Furthermore, we also categorize the languages involved in the downstream tasks into two groups based on the training data of PRM: seen languages (en, de, es, fr, ru, sw, and zh) and unseen languages (bn, ja, te, and th). To ensure the quality of our testset, we employ two human translators to post-edit the translated examples for each high-resource language (de, es, fr, ru, zh, and ja) and leverage gpt-4o to revise the translations in low-resource languages (bn, sw, te, and th). More details are shown in Appendix B.
MATH500 $\mu_{\textsc{all}}$ $\mu_{\textsc{seen}}$ $\mu_{\textsc{unseen}}$ en de es fr ru sw zh ja bn te th
MetaMath-Mistral-7B
Baseline 22.1 24.3 18.2 26.8 26.2 28.2 25.4 27.4 13.4 23.0 25.0 18.0 10.6 19.2
PRM-mono - 42.5 - 49.0 44.4 45.8 45.6 46.0 25.0 41.8 - - - -
PRM-cross 39.4 43.1 39.1 49.0 45.4 45.0 46.8 46.4 25.2 43.8 43.6 31.4 22.0 34.6
PRM-multi 39.6 43.1 39.4 50.2 45.6 47.4 45.4 45.2 25.2 42.8 43.6 32.6 21.8 35.2
Llama-3.1-8B-math
Baseline 22.1 24.3 18.1 30.4 22.4 27.4 25.4 22.0 15.4 27.4 20.0 16.6 16.0 19.8
PRM-mono - 43.3 - 49.0 46.2 45.8 44.2 45.8 26.2 46.2 - - - -
PRM-cross 40.9 43.6 36.3 49.0 48.8 46.6 44.8 44.8 26.0 45.2 43.0 36.0 28.2 37.8
PRM-multi 41.7 44.8 36.4 51.0 48.8 45.8 46.0 46.2 28.4 47.2 42.0 34.6 30.2 38.6
DeepSeekMath-7B-Instruct
Baseline 26.4 32.5 15.7 42.0 35.6 36.4 35.0 36.4 9.6 32.4 33.2 9.8 4.6 15.2
PRM-mono - 55.1 - 63.0 59.0 60.4 59.0 60.2 29.2 55.0 - - - -
PRM-cross 50.2 54.9 41.9 62.4 60.0 59.8 61.4 57.4 29.4 54.0 54.4 38.2 32.4 42.6
PRM-multi 51.3 55.6 43.7 63.8 58.6 60.2 60.2 61.4 30.6 54.2 55.8 38.0 35.6 45.4
Table 1: Different PRMs’ best-of-N sampling (N = 64) performance on MATH500 with the generator of MetaMath-Mistral-7B, Llama-3.1-8B-math, and DeepSeekMath-7B-Instruct. $\mu_{\textsc{all}}$, $\mu_{\textsc{seen}}$, and $\mu_{\textsc{unseen}}$ indicate the macro-average of results across all the languages, the seen languages, and the unseen languages, respectively.
Multilingual PRM Setups
To better understand PRMs in the context of multilingual research, we define three setups: PRM-mono, PRM-cross, and PRM-multi. The PRM-mono setup is trained and evaluated on the same single language, serving as the baseline for monolingual PRMs. The PRM-cross setup is trained on one language but evaluated on all 11 test languages. Specifically, in this work, we train PRM-cross on the English PRM dataset unless otherwise specified. Finally, the PRM-multi setup represents the multilingual PRM, which is both trained on all the seen languages and evaluated on all 11 test languages. To enhance the reliability and generalizability of our study, we train our multilingual PRM (verifier) based on the Qwen2.5-Math-7B-Instruct (Yang et al., 2024a), and leverage three diverse LLMs as the generator: MetaMath-Mistral-7B (Yu et al., 2023b), Llama-3.1-8B-math (fine-tuned with the MetaMath dataset (Dubey et al., 2024)), [^2] https://huggingface.co/gohsyi/Meta-Llama-3.1-8B-sft-metamath and DeepSeekMath-7B-Instruct (Shao et al., 2024). The details of training these PRMs are presented in Appendix C.
5 Recipes for Multilingual PRM Training
In this section, we conduct a series of experiments to investigate the performance of multilingual PRM. We examine how PRM-multi compares to PRM-mono and PRM-cross (Section 5.1), the impact of the number of training languages (Section 5.2), and the effect of varying the proportion of English in the training data (Section 5.3).
5.1 Monolingual, Cross-lingual, or Multilingual PRMs?
Building on Wu et al. (2024b)’s findings that cross-lingual ORMs outperform monolingual ones, we investigate the impact of multilingualism on PRMs. Specifically, we compare PRM-mono, PRM-cross, and PRM-multi to determine which setup offers best performance across languages.
Setup
We include three setups in this work. The PRM-mono is trained and evaluated on each individual language from the set of seen languages. The PRM-cross is trained exclusively on an English dataset and evaluated on all 11 test languages. Finally, the PRM-multi is trained on all seen languages and tested on all 11 test languages.
Multilingual PRMs perform best, followed by cross-lingual PRMs, while monolingual PRMs achieve the worst performance, on the seen languages.
As shown in Table 1, PRM-multi consistently achieves the highest performance across multiple language generators on the seen languages, surpassing PRM-mono and PRM-cross by +1.5 and +1.2 with Llama-3.1-8B-math generator, respectively. This indicates that incorporating data from multiple languages for PRM training significantly enhances the model’s ability across different languages. When comparing PRM-mono and PRM-cross, we observe that PRM-cross outperforms the PRM-mono for the English-centric MetaMath-Mistral-7B and Llama-3.1-8B-math generators. We hypothesize that this advantage stems from the pre-training phase: these generators are predominantly trained on English data but have limited exposure to multilingual corpora. As a result, fine-tuning on English PRM data enhances the reasoning capabilities of PRMs, facilitating greater cross-lingual transfer. More monolingual results are in Appendix D.
Multilingual PRMs generalize better on the unseen languages.
Both PRM-cross and PRM-multi are evaluated on four additional unseen languages. As shown in Table 1, PRM-multi demonstrates superior overall performance on the unseen languages in terms of $\mu_{\textsc{unseen}}$. These results suggest that training PRMs on multilingual datasets can effectively enhance model generalization to the unseen languages. More results on general-purpose LLM are provided in Appendix F.
In conclusion, these findings demonstrate that training a single multilingual PRM is an effective strategy for broad cross-lingual coverage, outperforming models trained either on a target language or on English alone. This outcome supports that PRM-multi is particularly advantageous for expanding the capabilities of PRMs in multilingual settings. More results on MGSM are in Appendix E.
5.2 Does More Languages Lead to Better Multilingual PRMs?
Figure 2: Best-of-N Performance on MGSM of PRMs trained using various subsets of English, German, Spanish, French, Russian, Swahili, and Chinese, with the generator of Llama-3.1-8B-math. The averages scores across all 11 languages.
While multilingual PRMs have demonstrated significant improvements, the question of how many languages are needed to achieve the best performance remains an open research problem. In this section, we address this research question by exploring the relationship between the number of training languages and the resulting performance.
Setup
We conduct experiments by training PRMs on datasets ranging from a single language up to all seven languages.
In this section, the number of total training examples of all PRMs are fixed. When the number of languages exceeds one, the total training examples are evenly distributed across all the selected languages. For evaluation, we test all PRMs on 11 different languages. The evaluation scores are averaged for each test language across all PRMs trained with the same number of languages.
More languages do not result in better multilingual PRMs.
As shown in Figure 2, the overall performance (AVG) improves as the number of training languages increases up to five languages. Beyond this point, adding more languages does not lead to further gains. Additionally, results from five individual languages (four seen languages and one unseen language) demonstrate that, although the optimal number of training languages varies across these languages, increasing the number of languages never leads to better performance. These findings suggest that increasing the number of training languages does not necessarily enhance multilingual PRMs. A key reason for this is the fixed amount of training data: as the number of languages grows, the training examples per language decrease. This reduction hinders sufficient training for seen languages and negatively impacts cross-lingual transfer to unseen languages.
5.3 How Much English Data Do We Need for Multilingual PRMs?
Figure 3: Best-of-N sampling performance of Llama-3.1-8B-math with PRMs finetuned on a training set where P% of the data is in English and (100 - P)% is uniformly distributed across six other languages. Each tick on the X-axis represents a specific tuning set configuration. The dash lines in blue, red, and green, indicate the average scores of all the languages, the seen languages, and the unseen languages, respectively.
While multilingual training with equal number of training examples in each language (PRM-multi) generally improves performance compared to English-only training (PRM-cross), we observe some exceptions on certain languages, as shown in Table 1. This observation prompts us to investigate how varying the number of English examples can affect the multilingual PRMs.
Setup
To explore this, we create data mixtures with varying percentages of English examples ($P\%$), with the remaining $(100-P)\%$ examples evenly distributed among six languages: German, Spanish, French, Russian, Swahili, and Chinese. Each PRM trained on these mixtures is then evaluated across all 11 languages.
Moderate amount of English data can lead to better multilingual PRMs.
As shown in Figure 3, incorporating a small amount of English data into the training mixture can lead to notable performance improvements across languages. Specifically, even as little as 1% of English examples significantly enhances performance, particularly for unseen languages. Interestingly, the majority of performance gains occur when English data constitutes less than 50% of the training mixture. However, when the proportion of English data exceeds 50%, performance begins to decline slightly across languages. Furthermore, training on 70% English data outperforms training solely on English (100%), suggesting that retaining some multilingual data introduces valuable variation and enhances the generalization capacity of multilingual PRMs. These findings indicate that as the proportion of English data increases, the PRMs may not be adequately trained on other seen languages, and unseen languages may benefit less from cross-lingual transfer. This highlights the importance of maintaining diverse and balanced language representation in multilingual training for optimal performance.
6 Analysis
In this section, we present a comprehensive analysis of our multilingual PRM, focusing on five critical aspects: error positions (Section 6.1), number of solutions (Section 6.2), integration of LoRA with PRM (Section 6.3), comparative evaluation with multilingual ORM (see Section 6.4), and implement PPO with multilingual PRM (see Section 6.5).
6.1 Which Steps Are More Prone to Errors?
Figure 4: Percentage distribution of the first error positions corresponding to the step in the reasoning on the PRM800K testset.
Figure 5: Best-of-N sampling performance of Llama-3.1-8B-math using different verification strategies across distinct numbers of solutions on MATH500.
PRMs provide fine-grained feedback on each intermediate step of a model’s chain-of-thought reasoning process. Errors at intermediate steps can propagate through the reasoning chain, ultimately affecting the final answer. Therefore, we investigate the earliest errors made by PRMs during the reasoning process, following Zheng et al. (2024).
MATH500
Mistral Llama DeepSeek
Verifier $\mu_{\textsc{all}}$ $\mu_{\textsc{seen}}$ $\mu_{\textsc{unseen}}$ $\mu_{\textsc{all}}$ $\mu_{\textsc{seen}}$ $\mu_{\textsc{unseen}}$ $\mu_{\textsc{all}}$ $\mu_{\textsc{seen}}$ $\mu_{\textsc{unseen}}$
Baseline 22.11 24.34 18.20 22.07 24.34 18.10 26.38 32.48 15.70
SC 29.20 31.80 24.65 30.60 33.31 25.85 44.96 49.29 37.40
ORM 39.54 42.63 34.25 40.49 43.14 35.85 50.96 55.54 42.95
PRM-multi 39.55 43.11 33.30 41.71 44.77 36.35 51.25 55.57 43.70
MGSM
Mistral Llama DeepSeek
Verifier $\mu_{\textsc{all}}$ $\mu_{\textsc{seen}}$ $\mu_{\textsc{unseen}}$ $\mu_{\textsc{all}}$ $\mu_{\textsc{seen}}$ $\mu_{\textsc{unseen}}$ $\mu_{\textsc{all}}$ $\mu_{\textsc{seen}}$ $\mu_{\textsc{unseen}}$
Baseline 49.63 61.65 28.60 56.18 64.23 42.10 52.95 63.02 35.30
SC 56.51 69.37 34.00 63.13 74.57 43.10 70.76 75.37 62.70
ORM 64.84 76.40 44.60 65.20 77.43 43.80 74.44 79.00 66.45
PRM-multi 65.45 77.09 45.10 71.93 82.00 54.30 75.42 80.51 66.50
Table 2: Multilingual PRMs’ best-of-N (N = 64) sampling performance on MATH500 and MGSM with three generators: MetaMath-Mistral-7B, Llama-3.1-8B-math, and DeepSeekMath-7B-Instruct. We use Qwen2.5-Math-7B-Instruct to finetune the ORM and PRM-multi. $\mu_{\textsc{all}}$, $\mu_{\textsc{seen}}$, and $\mu_{\textsc{unseen}}$ indicate the macro-average of results across all the languages, the seen languages, and the unseen languages, respectively.
Setup
We select a subset of instances from the PRM800K Russian test set where the final answers made by PRM-mono, PRM-cross, and PRM-multi are incorrect. For these instances, we identify the first occurrences of incorrect predictions from these PRMs. We classify the first error positions into three groups: early (steps 1 to 5), middle (steps 6 to 10), and later (steps 11 to 15).
Multilingual PRMs produce fewer errors at early steps.
The distribution of the earliest error positions, visualized in Figure 4, reveals a clear distinction between the three PRM configurations. In both PRM-mono and PRM-cross, a significant proportion of errors occurs within the early steps. In contrast, PRM-multi demonstrates fewer errors within this range and exhibits a slightly higher number of errors in later steps. These observations suggest that PRM-multi may be less prone to error propagation in the reasoning process, enabling it to maintain a more reliable reasoning trajectory. Consequently, PRM-multi can effectively achieve better overall performance.
6.2 Do More Candidates Drive Better Performance?
Recent research suggests that providing more candidate solutions can significantly boost the performance of PRM (Wang et al., 2024a, b). To explore if this applies in multilingual settings, we examine the impact of varying the number of candidates on PRM-mono, PRM-cross, and PRM-multi.
Setup
We conduct experiments on the MATH500 benchmark using the Llama-3.1-8B-math generator to compare the performance of PRM-multi, PRM-cross, and PRM-mono. For each approach, we vary the number of candidates N from 2 to 64. This allows us to assess how the number of candidate solutions influences performance across different PRM strategies in a multilingual context.
Multilingual PRMs yield better performance with more candidate solutions.
Figure 5 illustrates that PRM-multi consistently outperforms both PRM-cross and PRM-mono, with its advantage growing more pronounced as the number of candidates (N) increases. This finding underscores the scalability of multilingual PRM in diverse linguistic scenarios. Overall, these observations reinforce the conclusion that multilingual PRM not only maintains superior performance but also scales well as more candidates are introduced.
6.3 Are Multilingual PRMs Compatible with Parameter-Efficient Finetuning?
Recent research has demonstrated the effectiveness of parameter-efficient finetuning (PEFT) across a variety of tasks (Houlsby et al., 2019; Li and Liang, 2021). Therefore, we explore whether the PEFT approaches, such as LoRA (Hu et al., 2022), also perform well on multilingual PRMs.
Setup
To investigate this question, we employ LoRA on the key, query, and value attention matrices. Specifically, we use a rank of 8 and a dropout rate of 0.05 for both multilingual and cross-lingual PRMs. We train for three epochs with a batch size of 64 and a learning rate of $1e^{-5}$.
LoRA is computationally efficient, but not as good as its fully-finetuning counterpart in multilingual PRMs.
Figure 6 demonstrates that fully fine-tuning (FFT) consistently outperforms LoRA in both cross-lingual and multilingual settings. The performance gap becomes larger on the MATH500 dataset, which contains more complex questions compared to MGSM, suggesting that FFT is better suited for tasks requiring deeper reasoning and understanding. These findings align with prior research, which indicates that while PEFT methods may fall short of FFT when tasks demand higher complexity or reasoning capabilities (Biderman et al., 2024). Interestingly, although LoRA-based methods generally lag behind FFT, multilingual LoRA achieves stronger results than cross-lingual LoRA. This highlights the benefits of leveraging multilingual data during parameter-efficient fine-tuning, as multilingual data likely provides richer data diversity and linguistic coverage.
Figure 6: Comparison between parameter-efficient finetuning (LoRA) PRM and fully fine-tuning (FFT) PRM with Llama-3.1-8B-math generator.
6.4 Does PRM Surpass ORM in the Multilingual Scenario?
In this section, we explore whether PRM also outperforms Outcome Reward Model (ORM) and self-consistency (SC) in multilingual settings.
Setup
Following Lightman et al. (2024); Wang et al. (2024b), we evaluate the performance of PRM-multi by comparing it with other verifier methods, including: Direct prediction (Baseline), Self-consistency (majority voting) (SC), and ORM. Specifically, we train a multilingual ORM using uniform example budgets across seven seen languages. Then we assess the performance of verifiers on seven seen languages as well as on four additional unseen languages.
Multilingual PRM outperforms SC and ORM across all languages and generators.
The results presented in Table 2 confirm that PRM consistently achieves higher accuracy on two benchmarks across multiple languages. Specifically, when using the Llama-3.1-8B-math as the generator, PRM improves average accuracy by +19.64 points on the MATH500 dataset and by +15.75 points on the MGSM dataset in terms of $\mu_{\textsc{all}}$, compared to the Baseline of direct prediction. These substantial gains suggest PRM’s potential to enhance reasoning performance in a multilingual setting. Furthermore, PRM also surpasses both SC and ORM. For example, PRM exceeds SC and ORM by margins of up to +8.80 and +6.73 points on MGSM, respectively, when using Llama-3.1-8B-math as the generator. Additionally, PRM demonstrates performance improvements for both seen and unseen languages. With the DeepSeekMath-7B-Instruct generator on MGSM, PRM achieves respective gains of +17.49 and +31.20 for the seen and unseen language sets, compared to the Baseline.
6.5 Can Multilingual PRM Enhance LLMs?
Baseline PPO-ORM PPO-PRM
English 78.40 80.40 82.40
German 68.80 64.00 68.80
Spanish 72.00 71.20 76.00
French 67.60 68.00 71.60
Russian 69.60 68.40 71.20
Swahili 33.60 38.80 41.20
Chinese 59.60 64.00 62.80
Japanese 48.80 46.80 49.20
Bengali 45.20 41.20 40.40
Telugu 17.60 20.40 18.00
Thai 56.80 51.20 56.80
Average 56.18 55.85 58.04
Table 3: Zero-shot evaluation on MGSM for Llama-3.1-8B-math improved via PPO with PRM-multi.
In this section, we demonstrate that the multilingual PRM can be used as the reward model for finetuning the LLMs under a RL paradigm.
Setup
We design experiments to improve Llama-3.1-8B-math using RL where we adopt the PPO strategy (Schulman et al., 2017) on the MetaMathQA training set (Yu et al., 2023b). We then evaluate the resulting policy models on MGSM using top-1 accuracy in a zero-shot setting. Due to the computational constraints, we only generate one response during the fine-tuning process.
Reinforcement learning with multilingual PRM further improves the performance of LLMs.
The results shown in Table 3 indicate that step-by-step PPO with PRM-multi (PPO-PRM) consistently outperforms a standard supervised fine-tuned Baseline and PPO with ORM (PPO-ORM). Llama-3.1-8B-math with PPO-PRM achieves average boosts of +1.86 and +2.19 across 11 languages, compared to Baseline and PPO-ORM, respectively. These findings highlight the importance of fine-grained multilingual rewards. These gains demonstrate that process rewards can refine policy decisions for both reasoning steps and final outputs with RL. More results are in Appendix G.
7 Conclusion
Through comprehensive evaluations spanning 11 languages, our work demonstrates that multilingual PRMs significantly enhance the ability to perform complex, multi-step reasoning tasks in various languages, consistently outperforming both monolingual and cross-lingual counterparts. Furthermore, our findings highlight that PRM performance is sensitive to the number of languages and the volume of English training data. The multilingal PRMs also benefit from more candidate responses and model parameters. These results underscore the importance of diverse language training in providing fine-grained rewards and open up promising avenues for multilingual reasoning.
8 Limitations
While we have demonstrated the effectiveness of multilingual PRMs, our study has not comprehensively explored the wide range of reward optimization methods (Rafailov et al., 2024; Azar et al., 2024), some of which may not benefit from cross-lingual reward model transfer. Nevertheless, best-of-N and PPO, the two techniques leveraged in this paper, are highly representative of current practices, particularly given the consistently strong performance of best-of-N (Gao et al., 2023; Rafailov et al., 2024; Mudgal et al., 2023). Furthermore, while our results show that multilingual PRMs outperform both cross-lingual and monolingual PRMs, our experiments are limited to 11 languages. Extending this approach to a broader set of languages and evaluating its impact across diverse linguistic families is an important avenue for future work.
Acknowledgement
This work is funded by EU Horizon Europe (HE) Research and Innovation programme grant No 101070631, and UK Research and Innovation under the UK HE funding grant No 10039436.
The computations described in this research were performed using the Baskerville Tier 2 HPC service (https://www.baskerville.ac.uk/). Baskerville was funded by the EPSRC and UKRI through the World Class Labs scheme (EP/T022221/1) and the Digital Research Infrastructure programme (EP/W032244/1) and is operated by Advanced Research Computing at the University of Birmingham.
Appendix A Data Statistics
#exam. max min mean
PRM800K trainset 404K 56 1 06.39
Math-Shepherd 445K 30 1 06.23
PRM800K testset 5071 53 1 22.11
Table 4: Dataset statistics of the datasets in this work, including number of examples, maximum, minimum, and average number of steps in the answers.
The dataset statistics are summarized in Table 4. These include the total number of examples, as well as the maximum, minimum, and average number of reasoning steps in the answers across all examples. For the selection criteria for the six languages, there are two key desiderata for the language selection in our work. Firstly, the examples must be accurately translatable into the target language by MT systems. Secondly, the target language must allow for proper evaluation. With these desiderata in mind, we selected six high-resource languages covered by the MGSM dataset. This choice ensures that the translated data closely aligns with the original English dataset and allows us to focus on comparing model strategies without introducing the added variability that lower-resource language translations might cause. We will clarify this in our future revision.
Appendix B Translation Details
Due to imbalanced resources across languages, translation has become a standard method for multilingual research. Recent research has demonstrated that machine-translated datasets are comparable to human-translated ones and can be directly used for training and evaluation (Chen et al., 2024; Thellmann et al., 2024).
In this study, after translating the English dataset into foreign languages, we use regular expressions to filter out the translated training instances that contain discrepancies in numbers or equations compared to the original English dataset. This ensures the correctness of the mathematical content. For the translated multilingual MATH500 test set, we employ two human translators to post-edit the test instances in high-resource languages (de, es, fr, ru, zh, and ja) by correcting inaccurate translations and verifying the consistency of mathematical notations. We pay \$0.05 USD for each example, resulting in a total cost of \$150 USD for post-editing. For the low-resource languages (bn, sw, te, and th) in MATH500, we leverage gpt-4o to post-edit the translations.
To verify the quality of our translations, we use Google Translate to back-translate the multilingual MATH500 and 1,000 random training instances from each training set into English. We then calculate the BLEU score using the original English instances as the reference translation. As shown in Table 5, the high BLEU scores confirm the quality of the translations in our datasets.
de es fr ru sw zh ja bn te th
Train 81.2 88.4 87.3 74.0 87.3 80.8 - - - -
Test 85.9 91.5 91.0 73.0 90.3 84.4 84.3 65.3 65.8 80.9
Table 5: BLEU scores of back-translation examples with using the original English data.
Appendix C Training Details
We train the PRMs by fine-tuning all parameters of Qwen2.5-Math-7B-Instruct using the AdamW optimizer with a learning rate of $10^{-5}$ and a batch size of 8. This process is conducted over two epochs on 4 NVIDIA A100 GPUs (80GB). During training, we use a linear learning rate schedule with a warm-up phase that constitutes 10% of the total training steps.
Appendix D Cross-lingual Transfer of PRMs
Figure 7: Performance of PRM-mono trained on seven seen languages and evaluated on all 11 languages based on the MATH500 with Llama-3.1-8B-math generator.
Following Wu et al. (2024b), we assess the performance of cross-lingual PRMs to inspect if language similarity like the script or mutual intelligibility might affect the levels of reasoning verification cross-lingual transfer.
Setup
We train PRMs on monolingual versions of the data in German, Spanish, French, Russian, Swahili, and Chinese, and evaluate their transfer to other languages.
No clear signal indicates that language similarity strongly correlates with cross-lingual transfer.
We present the cross-lingual transfer results in Figure 7 and observe that there is no clear conclusion regarding the factors that impact cross-lingual transfer. For instance, the PRM trained on Russian data achieves the highest accuracy when evaluating French, Swahili, Chinese, Telugu, and Thai. Notably, these languages neither share the same script nor belong to the same language family as Russian. This observation suggests that linguistic similarity, in terms of script or language family, may not be a decisive factor in cross-lingual transfer. These findings underscore the uncertainty in predicting cross-lingual transfer performance based solely on language similarity. In practice, selecting a diverse set of representative languages for training a multilingual PRM may be a more effective strategy to address this uncertainty and improve performance across a wide range of target languages.
Appendix E Breakdown Results of MGSM for PRM-mono, PRM-cross, and PRM-multi
MGSM $\mu_{\textsc{all}}$ $\mu_{\textsc{seen}}$ $\mu_{\textsc{unseen}}$ en de es fr ru sw zh ja bn te th
MetaMath-Mistral-7B
PRM-mono - 76.0 - 90.8 78.8 81.2 81.6 86.0 36.0 77.6 - - - -
PRM-cross 65.2 76.7 45.2 90.8 84.4 85.2 82.4 86.8 27.2 80.0 76.2 43.0 7.6 54.0
PRM-multi 65.5 77.1 45.1 89.2 83.2 86.0 82.4 86.4 33.2 79.2 75.6 43.2 8.0 53.6
Llama-3.1-8B-math
PRM-mono - 81.7 - 92.4 83.2 88.0 80.4 82.4 62.4 83.2 - - - -
PRM-cross 68.8 79.3 50.6 92.4 82.0 88.0 82.0 79.2 50.4 80.8 72.8 39.6 20.8 69.2
PRM-multi 71.9 82.0 54.3 90.4 87.6 88.0 83.6 83.2 59.6 81.6 74.0 48.0 23.6 71.6
DeepSeekMath-7B-Instruct
PRM-mono - 80.5 - 96.4 86.4 90.4 85.2 88.0 32.0 85.0 - - - -
PRM-cross 74.0 79.0 65.1 96.4 86.0 91.2 85.6 87.2 18.4 88.4 80.0 57.6 51.6 71.2
PRM-multi 75.4 80.5 66.5 95.2 84.0 92.4 86.4 89.2 30.0 86.4 80.8 60.8 52.4 72.0
Table 6: Different PRMs’ best-of-N sampling (N = 64) performance on MGSM with the generator of MetaMath-Mistral-7B, Llama-3.1-8B-math, and DeepSeekMath-7B-Instruct. $\mu_{\textsc{all}}$, $\mu_{\textsc{seen}}$, and $\mu_{\textsc{unseen}}$ indicate the macro-average of results across all the languages, the seen languages, and the unseen languages, respectively.
We present the breakdown of results for each language on the MGSM in Table 6. The results indicate that the PRM-multi consistently outperforms both the PRM-mono and PRM-cross models across languages. This observation aligns with the conclusion drawn in Section 5.1, highlighting the advantages of multilingual training for PRMs.
Appendix F Results on General-Purpose LLM
We provide the results of the general LLM Qwen2.5-7B-Instruct on MATH500 in Table 7. It can be observed that the multilingual PRM achieves consistent conclusions when applied to the general LLM.
MATH500 $\mu_{\textsc{all}}$ $\mu_{\textsc{seen}}$ $\mu_{\textsc{unseen}}$ en de es fr ru sw zh ja bn te th
Baseline 36.3 37.9 33.4 44.2 40.6 41.6 40.2 40.8 20.8 37.4 40.6 38.2 20.2 34.4
PRM-mono - 54.2 - 61.0 57.6 58.0 57.0 58.2 31.6 56.2 - - -
PRM-cross 53.2 57.1 55.7 63.6 60.8 60.4 60.4 61.6 33.8 59.4 60.8 58.8 36.8 56.4
PRM-multi 54.0 58.2 56.7 64.8 61.6 61.8 61.2 62.2 35.2 60.6 61.2 58.6 38.2 57.8
Table 7: Performance on general LLM Qwen2.5-7B-Instruct.
Checkpoint Bengali English French
Baseline 45.2 78.4 67.6
Checkpoint-500 46.8 80.0 69.2
Checkpoint-1150 (final) 40.4 82.4 71.6
Table 8: The influence of final checkpoint selection strategy during the PPO training process.
Appendix G Influence of Checkpoint Selection
We observe a decline in Bengali performance in both ORM and PRM, as shown in Table 3. Upon evaluating the performance at each intermediate checkpoint, our analysis indicates that this behavior stems from the PPO training process and the strategy used for selecting the final checkpoint, as illustrated in Table 8. Specifically, since the checkpoint is selected based on the average loss across all languages, the one that minimizes the overall loss does not necessarily yield optimal performance for individual languages. In this case, Bengali appears to follow a distinct learning rate trajectory compared to other languages. We acknowledge this limitation and plan to investigate language-specific adjustments to the training process in future work.